OpenStudy (anonymous):

5. Which of the following operation are commutative? division and subtraction multiplication and addition multiplication and division addition and subtraction

OpenStudy (anonymous):

hi. are you still stuck?

OpenStudy (anonymous):

multiplication and addition

OpenStudy (anonymous):

3*2 = 2*3 3+2 = 2+3

OpenStudy (anonymous):

coolsector is correct. these are properties of real numbers that come in handy in solving equations.

OpenStudy (anonymous):

thanks